Conduct a traditional hypothesis test for the null hypothesis H0: µ = 100 against the alternative


hypothesis H1: µ > 100 based on the 35 random observations. The sample mean is 105 and the


sample standard deviation is 15. Use α = 0.01.

1
Expert's answer
2022-05-30T23:33:57-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=100"

"H_1:\mu>100"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=34" and the critical value for a right-tailed test is "t_c =2.44115."

The rejection region for this right-tailed test is "R = \{t:t>2.44115\}."


The t-statistic is computed as follows:



"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{105-100}{15/\sqrt{35}}=1.9720"


Since it is observed that "t=1.9720<2.44115=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for right-tailed, "df=34" degrees of freedom, "t=1.9720" is "p= 0.028395," and since "p= 0.028395>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is greater than 100, at the "\alpha = 0.01" significance level.